Please help me solve this equation:
Miki had four times as much money as Joanne. When grandma gave each $20, Miki had twice as much as Joanne. How much did each have originally?

Answer by FrankM(1040)

M=4J
M+20=2(J+20)
M+20=2J+40
M=2J+20
4J=2J+20
2J=20
J=10
M=40
when you add 20, it's 30 and 60